# Enhanced thermoelectric effects in a driven one-dimensional system

### v1

@@ IV.1.1 Thermoelectric effects of AC driving
We can now calculate how the thermoelectric properties of the system are affected by the periodic driving. We consider the Seebeck coefficient $S$, which is derived in Sec. II in Eq. (6). As discussed in Sec. II, in the low-temperature limit ($k_{B}T\ll\mu$) $S$ grows linearly with temperature $T$. In order to emphasize the deviation from the low-temperature behavior of the Seebeck coefficient, in Fig. 3 we plot the quantity $S/T\times(\mu/k_{B})$ as a function of $k_{B}T/\mu$ for a fixed amplitude of oscillations $v=0.4$ and different values of frequency: $\hbar\omega/\mu=0.2$ (blue curve), $0.5$ (red curve), $0.8$ (purple), $0.9$ (green curve). The black dashed curve represents the static situation ($v=0$), plotted for reference.
@@
The black dashed curve is constant for small values of $T$, as expected, but above $k_{B}T=0.1\mu$ starts to increases roughly linearly with temperatures (i.e. $S\propto T^{2}$). The blue curve in Fig. 3 has a similar behavior to the black dashed curve except for an enhancement at low-temperatures due to the pronounced kink structure in the transmission [see the solid blue curve in Fig. 2(a)]. Specifically, the kink at $E=\hbar\omega$ leads to a corresponding suppression of $\mathcal{T}(\mu)$ of about 1%, which appears in the denominator of Eq. (9). Conversely, $\mathcal{T}^{\prime}(\mu)$, present in the numerator of Eq. (9), experiences a much larger increase of about 30%. This phenomenon is illustrated in Fig. 2(a) by comparing the blue and black solid curves. An even more pronounced effect is observed when the frequency of the AC driving increases, so that the kink structure approaches the chemical potential. Indeed, in Fig. 3, we observe an increase of $S/T$ at low temperatures for increasing AC frequencies. Furthermore, at zero temperature, one observes a small peak in the red curve ($\hbar\omega=0.5\mu$). This peak is an effect of the second harmonic of the AC driving (two photon process), which slightly changes the transmission function around the chemical potential (i.e.  at $E=\mu$) as discussed above. Moreover, this shows that periodic driving could potentially increase the thermoelectricity when the driving frequency satisfies resonant conditions $\hbar\omega\simeq\mu/n$, with $n$ an integer number.
@@
We note that the green curve in Fig. 3 shows a significantly larger value of $S/T$ at low temperatures compared to the static case (black dashed curve). However, a strikingly different behavior emerges at higher temperatures, specifically above $k_{B}T\simeq 0.02\mu$, where $S/T$ decreases dramatically. This decline occurs because, in the thermal window around the chemical potential, the suppression of the transmission due to the kink located around $E-\mu=-0.1\mu$ starts to have an effect. This behavior is discussed in details in App. C. Finally, Fig. 3 shows that in the high-temperature regime (when $\hbar\omega\lesssim k_{B}T$) the AC driving only slightly corrects the static regime.
@@
Figure 4: Zero-temperature limit of the normalized Seebeck coefficient, $S/T\times(\mu/k_{B})$, in units of $k_{B}/e$, as a function of $\hbar\omega/\mu$. (a) For different $\eta$ values: $\eta=0.1$ (red), $0.2$ (orange), $0.3$ (green), $0.4$ (black) and a fixed strength of oscillating delta-barrier $v=0.4$. (b) For different $v$ values: $v=0.1$ (purple), $0.2$ (red), $0.3$ (blue), $0.4$ (green) and a fixed $\eta=0.3$. The dash lines are $\mathcal{T}^{\prime}(\mu)/\mathcal{T}(\mu)$ multiplied by $\pi^{2}/3$ as reference. The colorful full circles are the same limit for static $v=0$ case.
@@
We can now examine how the parameters $v$ and $\eta$ influence the behavior of the Seebeck coefficient $S$. To this end, we present in Fig. 4 the ratio $S/T$, calculated in the zero-temperature limit, as a function of the normalized frequency $\hbar\omega/\mu$ for various values of $v$ and $\eta$. For completeness, we also report, with colored circles, the value of $S/T$ in zero-temperature limit relative to the static case ($v=0$). It is important to note that, indeed, the small frequency limit, where $\omega\to 0$ and $v$ has a finite fixed value, does not necessarily coincide with the static case computed for $v=0$. The non-commutativity of the limits $\omega\to 0$ and $v\to 0$ reflects the resonant, non-perturbative nature of the photon-assisted processes. These non-perturbative effects are the origin of the kink structures observed in the transmission functions.
@@
In Fig. 4(b), we systematically investigate the low temperature limit of the ratio $S/T$ as a function of frequency fixing the strength of the static component of the potential to $\eta=0.3$. We plot four different curves corresponding to different values of the oscillating amplitude $v$ (see caption). It shows an overall, though not pronounced, increase of $S/T$ with increasing $v$. More precisely, the downward bend occurring for $\hbar\omega>\mu/2$, as well as the sharp features at $\hbar\omega=\mu/2$, become more evident by increasing $v$. Furthermore, the effect of the third harmonic, at $\hbar\omega=\mu/3$, becomes observable only for the largest value of $v=0.4$ considered (green curve).
@@
where $S_{0}$ represents the Seebeck coefficient under static conditions (i.e., when $v=0$). In Fig. 5 we plot the quantity $R_{S}$ as a function of temperature for four values of frequency: $\hbar\omega/\mu=0.2$ (blue), $0.5$ (red), $0.8$ (purple), $0.9$ (green) for the parameters which yield the largest variation, namely $v=0.4$ and $\eta=0.1$. All curves show an enhancement of the Seebeck coefficient at low temperatures. The blue curve ($\hbar\omega=0.2\mu$) shows an increment of about 30% in a temperature region up to $T=0.1\mu/k_{B}$. As the frequency increases, the kink in the transmission approaches the chemical potential causing the plateau in $R_{S}(T)$ to become higher but narrower. For the red curve ($\hbar\omega=0.5\mu$) the enhancement is about 40%, in the temperature range from 00 to $0.08\mu/k_{B}$. The zero-temperature peak is due to the double-frequency effect, as mentioned before. For the purple curve ($\hbar\omega=0.8\mu$) the enhancement is about 70%, in the temperature range from 00 to $0.05\mu/k_{B}$, while for $\hbar\omega=0.9\mu$ (green curve) the enhancement reaches 200%, but in a smaller temperature region from 00 to $0.02\mu/k_{B}$.
In conclusion, for system with quite low electron density (i.e., small Fermi energy) the AC driving could indeed increase the Seebeck coefficient in the quantum regime $k_{B}T\ll\hbar\omega\lesssim\mu$.
@@ IV.2 Single barrier with a step
In this section, we consider the scenario with a step function potential and set $\Delta>0$ in the Hamiltonian given by Eq. (10). Since $\Delta$ breaks spatial inversion symmetry the transmission functions from left to right $\mathcal{T}_{RL}$ and from right to left $\mathcal{T}_{LR}$ are not necessarily equal when an AC driving is applied.
@@
Figure 6: Single barrier with a step potential: energy dependence of transmission functions. (a) Blue and green curves are for the left-to-right transmission $\mathcal{T}_{RL}$, while red and magenta curves are for right-to-left transmission $\mathcal{T}_{LR}$. One set of curves (red and blue) refer to the step height $\Delta=0.75\mu$, and the other (green and magenta) refer to $\Delta=1.1\mu$. The inset is the zoom-in of the kink structures around $E=0.95\mu$. (b) Blue (green) curve is the average transmission $\bar{\mathcal{T}}=(\mathcal{T}_{RL}+\mathcal{T}_{LR})/2$ for $\Delta=0.75\mu$ ($\Delta=1.1\mu$). For both panels, black dashed curves refer to the static case with $\Delta=0.75\mu$, while black dash-dotted curves correspond to the static case with $\Delta=1.1\mu$. Vertical dotted lines are guides for the eyes, indicating the corresponding frequencies of the AC signal. Vertical thin gray lines indicate the chemical potential, i.e. $E=\mu$. Other parameters are: $\eta=0.2$, $v=0.4$ and $\hbar\omega=0.2\mu$.
@@
In Fig. 6(a) we show these transmissions as functions of energy $E$ for a fixed frequency $\hbar\omega=0.2\mu$. The plot is divided into two transmission ranges: 0.5 to 0.9 (top panel) and 0 to 0.05 (bottom panel), allowing for a clearer visualization of features present at low transmission values. Two sets of curves illustrate two distinct scenarios with different step heights: for $\Delta=0.75\mu$ (blue and red curves) the chemical potential is above the step height, while for $\Delta=1.1\mu$ (green and magenta curves) the chemical potential is below the step height. The blue and green curves represent the transmissions $\mathcal{T}_{\rm RL}$, while the red and magenta curves correspond to $\mathcal{T}_{\rm LR}$. Dashed curves depict the static case, included for reference, for which the equality $\mathcal{T}_{RL}=\mathcal{T}_{LR}$ holds. In the static case, as illustrated in Fig. 6(a), transmission is zero for energies below the step ($E<\Delta$), since the electronic states on the right lead are evanescent (do not propagate). In contrast, AC driving leads to a non-zero transmission $\mathcal{T}_{RL}$ even for energies below $\Delta$, as evident from the green and blue curves in the figure. This is due to the fact that an incident electron from the left lead, possessing energy $E<\Delta$, can overcome the potential step and transfer to the right lead into a higher energy state ($E+n\hbar\omega>\Delta$) by absorbing $n$ energy quanta through AC driving. On the other hand, for $E<\Delta$, the red and magenta curves (right-to-left transmission, $\mathcal{T}_{LR}$) exhibit behavior analogous to the static case: they are zero because incident states from the right lead cannot propagate below the step. Figure 6(a) also shows that for energies above the step, all AC-driven curves lie slightly below their static counterparts (dashed lines).
@@
Similar to the findings in Section IV.1, albeit less pronounced, additional kink structures appear at $E=\Delta+\hbar\omega$ (marked by the vertical dotted lines). To enhance visibility of these features, in Fig. 6(a) we include an inset showing a zoom-in of the red and blue curves ($\Delta=0.75\mu$). Note that the kinks for the two transmissions in the blue and red curves go in opposite directions (one goes down and the other goes up, respectively), so that when summed together they may even compensate. This is shown in Figure 6(b) where we plot the averaged transmissions $\bar{\mathcal{T}}=(\mathcal{T}_{RL}+\mathcal{T}_{LR})/2$. Note that here the range of values of transmissions is different from the one used in panel (a) and ranges between 0 and 0.9. Finally, both Figs. 6(a) and (b) show that there is no qualitative difference between the group of curves for $\Delta=1.1\mu$ and $\Delta=0.75\mu$, apart from the obvious shift in energy.
@@
Before proceeding, we note that because of the difference between $\mathcal{T}_{RL}(\epsilon)$ and $\mathcal{T}_{LR}(\epsilon)$, a non-zero pumping current can flow through the system even in the absence of voltage or thermal biases. Indeed, according to Eqs. (3) and (5), in this case the charge current is given by
[…]
@@
More details are reported in Appendix B. We can now study the thermoelectric properties of the system and calculate the coefficient $L_{\rm eh}$ defined in Eq. (8). Here we do not consider the Seebeck coefficient $S$ because the charge current, and hence the coefficient $L_{\rm ee}$, is strongly suppressed when the chemical potential $\mu$ is lower than the height of the step $\Delta$. Indeed, in this situation the behavior of $S$ is totally dominated by $L_{\rm ee}$, see Eq. (6).
@@
Since in the static situation $L_{\rm eh}$ scales as $T^{3}$ in the low-temperature regime, see Sec. II, in Fig. 7 we plot the ratio $L_{\rm eh}/T^{3}$ as a function of the rescaled temperature $k_{B}T/\mu$ for the two values of potential step height: blue curve for $\Delta=0.75\mu$ and green curve for $\Delta=1.1\mu$. The dashed black curves refer for the static case. The main difference between the two is that, at $T=0$, one takes a finite value ($\Delta=0.75\mu$), while the other becomes zero ($\Delta=1.1\mu$). Both curves exhibit a broaden peak at intermediate values of temperatures (between 0.05 and 0.1 $k_{B}T/\mu$). Figure 7(a) shows that the AC driving, in both cases, causes $L_{\rm eh}$ to slightly increase at small $T$ and slightly decrease at higher $T$. In both cases, though, the two curves do not depart much from the static case. This behavior is discussed in details in App. C.
@@
where $L_{\rm eh}^{(0)}$ denotes the static thermoelectric coefficient, i.e., the value of $L_{\rm eh}$ when $v=0$. The quantity $R_{L_{\rm eh}}$ is plotted in Fig. 7(b) as a function of the normalized temperature for the two different values of $\Delta$, for fixed values of $\omega$, $v$, and $\eta$. The blue and green curves represent the cases with the step value $\Delta=0.75\mu$ and $\Delta=1.1\mu$, respectively. Both curves show an enhancement of $L_{\rm eh}$ in the low-temperature regime. The blue curve has a maximum of about 5% increase, while the green curve shows a much larger increase, since it diverges for $T$ going to zero. The inset shows a logarithmic scale zoom-in in a range of small temperatures. This huge increase is due to the photon-assisted effect of the driving which enhances the thermoelectric coefficient $L_{\rm eh}$ in a range of temperatures where transport is otherwise suppressed by the presence of the potential step.

### v2

@@ IV.1.1 Thermoelectric effects of AC driving
We can now calculate how the thermoelectric properties of the system are affected by the periodic driving. We consider the Seebeck coefficient $S$, which is derived in Sec. II in Eq. (6). As discussed in Sec. II, in the low-temperature limit ($k_{B}T\ll\mu$) $S$ grows linearly with temperature $T$. In order to emphasize the deviation from the low-temperature behavior of the Seebeck coefficient, in Fig. 3 we plot the quantity $S/T\times(\mu/k_{B})$ as a function of $k_{B}T/\mu$ for a fixed amplitude of oscillations $v=0.4$ and different values of frequency: $\hbar\omega/\mu=0.2$ (blue curve), $0.5$ (red curve), $0.8$ (purple curve), $0.9$ (green curve). The black dashed curve represents the static situation ($v=0$), plotted for reference.
@@
The black dashed curve is constant for small values of $T$, as expected, but above $k_{B}T=0.1\mu$ starts to increases roughly linearly with temperatures (i.e. $S\propto T^{2}$). The blue curve in Fig. 3 has a similar behavior to the black dashed curve except for an enhancement at low temperatures due to the pronounced kink structure in the transmission [see the solid blue curve in Fig. 2(a)]. Specifically, the kink at $E=\hbar\omega$ leads to a corresponding suppression of $\mathcal{T}(\mu)$ of about 1%, which appears in the denominator of Eq. (9). Conversely, $\mathcal{T}^{\prime}(\mu)$, present in the numerator of Eq. (9), experiences a much larger increase of about 30%. This phenomenon is illustrated in Fig. 2(a) by comparing the blue and black solid curves. An even more pronounced effect is observed when the frequency of the AC driving increases, so that the kink structure approaches the chemical potential. Indeed, in Fig. 3, we observe an increase of $S/T$ at low temperatures for increasing AC frequencies. Furthermore, at zero temperature, one observes a small peak in the red curve ($\hbar\omega=0.5\mu$). This peak is an effect of the second harmonic of the AC driving (two-photon process), which slightly changes the transmission function around the chemical potential (i.e.  at $E=\mu$) as discussed above. Moreover, this indicates that periodic driving may potentially enhance thermoelectricity when the driving frequency satisfies resonant conditions $\hbar\omega\simeq\mu/n$, with $n$ an integer.
@@
We note that the green curve in Fig. 3 shows a significantly larger value of $S/T$ at low temperatures compared to the static case (black dashed curve). However, a strikingly different behavior emerges at higher temperatures, specifically above $k_{B}T\simeq 0.02\mu$, where $S/T$ decreases dramatically. This decline occurs because, within the thermal window around the chemical potential, the suppression of transmission due to the kink located at $E-\mu=-0.1\mu$ begins to take effect. This behavior is discussed in detail in App. C. Finally, Fig. 3 shows that in the high-temperature regime (when $\hbar\omega\lesssim k_{B}T$) the AC driving only slightly corrects the static regime.
@@
Figure 4: Zero-temperature limit of the normalized Seebeck coefficient, $S/T\times(\mu/k_{B})$, in units of $k_{B}/e$, as a function of $\hbar\omega/\mu$. (a) For different $\eta$ values: $\eta=0.1$ (red), $0.2$ (orange), $0.3$ (green), $0.4$ (black) and a fixed strength of oscillating delta-barrier $v=0.4$. (b) For different $v$ values: $v=0.1$ (purple), $0.2$ (red), $0.3$ (blue), $0.4$ (green) and a fixed $\eta=0.3$. The dotted curves are $\mathcal{T}^{\prime}(\mu)/\mathcal{T}(\mu)$ multiplied by $\pi^{2}/3$ as reference. The colorful full circles are the same limit for the static $v=0$ case.
@@
We can now examine how the parameters $v$ and $\eta$ influence the behavior of the Seebeck coefficient $S$. To this end, we present in Fig. 4 the ratio $S/T$, calculated in the zero-temperature limit, as a function of the normalized frequency $\hbar\omega/\mu$ for various values of $v$ and $\eta$. For completeness, we also report, with colored circles, the value of $S/T$ in the zero-temperature limit relative to the static case ($v=0$). It is important to note that, indeed, the small frequency limit, where $\omega\to 0$ and $v$ has a finite fixed value, does not necessarily coincide with the static case computed for $v=0$. The non-commutativity of the limits $\omega\to 0$ and $v\to 0$ reflects the resonant, non-perturbative nature of the photon-assisted processes. These non-perturbative effects are the origin of the kink structures observed in the transmission functions.
@@
In Fig. 4(b), we systematically investigate the low temperature limit of the ratio $S/T$ as a function of frequency, fixing the strength of the static component of the potential to $\eta=0.3$. We plot four different curves corresponding to different values of the oscillating amplitude $v$ (see caption). It shows an overall, though not pronounced, increase of $S/T$ with increasing $v$. More precisely, the downward bend occurring for $\hbar\omega>\mu/2$, as well as the sharp features at $\hbar\omega=\mu/2$, become more evident as $v$ increases. Furthermore, the effect of the third harmonic, at $\hbar\omega=\mu/3$, becomes observable only for the largest value of $v=0.4$ considered (green curve).
@@
where $S_{0}$ represents the Seebeck coefficient under static conditions (i.e., when $v=0$). In Fig. 5 we plot the quantity $R_{S}$ as a function of temperature for four values of frequency: $\hbar\omega/\mu=0.2$ (blue), $0.5$ (red), $0.8$ (purple), $0.9$ (green) for the parameters which yield the largest variation, namely $v=0.4$ and $\eta=0.1$. All curves show an enhancement of the Seebeck coefficient at low temperatures. The temperature dependence of $R_{\rm S}$ is in general quite mild (see, for example, blue and purple curves in Fig. 5): $R_{\rm S}$ remains roughly constant in a range of temperatures and then decreases smoothly to 0. Exceptions are the case $\hbar\omega=0.5\mu$ (red curve), where a peak appears at zero temperature, and the case $\hbar\omega=0.9\mu$ of very high frequency (green curve), where $R_{\rm S}$ decreases dramatically, reaching negative values. Such a temperature sensitivity can be understood from the dependence on temperature of the normalized Seebeck coefficient $S$, reported in Fig. 3 and discussed above in this subsection. In particular, the curves in Fig. 5 exhibit the following behavior. The blue curve ($\hbar\omega=0.2\mu$) shows an increment of about 30% in a temperature region up to $T=0.1\mu/k_{B}$. As the frequency increases, the kink in the transmission approaches the chemical potential, causing the plateau in $R_{S}(T)$ to become higher but narrower. For the red curve ($\hbar\omega=0.5\mu$) the enhancement is about 40%, in the temperature range from 0 to $0.08\mu/k_{B}$. The zero-temperature peak is due to the double-frequency effect, as mentioned before. For the purple curve ($\hbar\omega=0.8\mu$) the enhancement is about 70%, in the temperature range from 0 to $0.05\mu/k_{B}$, while for $\hbar\omega=0.9\mu$ (green curve) the enhancement reaches 200%, but in a smaller temperature region from 0 to $0.02\mu/k_{B}$. Especially in the latter case, there is a trade-off between a large increase in the Seebeck coefficient and the maximum operating temperature.
In conclusion, for systems with quite low electron density (i.e., small Fermi energy) the AC driving could indeed increase the Seebeck coefficient in the quantum regime $k_{B}T\ll\hbar\omega\lesssim\mu$. We expect the effect to appear in semiconducting nanostructures, such as nanowires, with a gate-tunable barrier [69], where thermoelectric measurements can also be performed at cryogenic temperatures [70, 71, 72, 73]. In this case, the typical values for the Seebeck coefficient are within reach of experimental parameters. For example, reading from Fig. 3, for $v=0.4$, $\eta=0.1$ and $\hbar\omega/\mu=0.5$ (red curve), one obtains $S/T=0.07k_{B}^{2}/(e\mu)$ for $k_{B}T/\mu=0.1$. By choosing $\mu=1$ meV, one finds $S=0.6$ $\mu V/$K at frequencies of hundreds of GHz.
@@ IV.2 Single barrier with a step
In this section, we consider the scenario with a step function potential and set $\Delta>0$ in the Hamiltonian given by Eq. (10). Since $\Delta$ breaks spatial inversion symmetry, the transmission functions from left to right $\mathcal{T}_{RL}$ and from right to left $\mathcal{T}_{LR}$ are not necessarily equal when an AC driving is applied.
@@
Figure 6: Single barrier with a step potential: energy dependence of transmission functions. (a) Blue and green curves are for the left-to-right transmission $\mathcal{T}_{RL}$, while red and magenta curves are for right-to-left transmission $\mathcal{T}_{LR}$. One set of curves (red and blue) refers to the step height $\Delta=0.75\mu$, and the other (green and magenta) refers to $\Delta=1.1\mu$. The inset is the zoom-in of the kink structures around $E=0.95\mu$. (b) Blue (green) curve is the average transmission $\bar{\mathcal{T}}=(\mathcal{T}_{RL}+\mathcal{T}_{LR})/2$ for $\Delta=0.75\mu$ ($\Delta=1.1\mu$). For both panels, black dashed curves refer to the static case with $\Delta=0.75\mu$, while black dash-dotted curves correspond to the static case with $\Delta=1.1\mu$. Vertical dotted lines are guides for the eyes, indicating the corresponding frequencies of the AC signal. Vertical thin gray lines indicate the chemical potential, i.e. $E=\mu$. Other parameters are: $\eta=0.2$, $v=0.4$ and $\hbar\omega=0.2\mu$.
@@
In Fig. 6(a), we show these transmissions as functions of energy $E$ for a fixed frequency $\hbar\omega=0.2\mu$. The plot is divided into two transmission ranges: 0.5 to 0.9 (top panel) and 0 to 0.05 (bottom panel), allowing for a clearer visualization of features present at low transmission values. Two sets of curves illustrate two distinct scenarios with different step heights: for $\Delta=0.75\mu$ (blue and red curves), the chemical potential is above the step height, while for $\Delta=1.1\mu$ (green and magenta curves), the chemical potential is below the step height. The blue and green curves represent the transmissions $\mathcal{T}_{\rm RL}$, while the red and magenta curves correspond to $\mathcal{T}_{\rm LR}$. Dashed black curves depict the static case, included for reference, for which the equality $\mathcal{T}_{RL}=\mathcal{T}_{LR}$ holds. In the static case, as illustrated in Fig. 6(a), transmission is zero for energies below the step ($E<\Delta$), since the electronic states on the right lead are evanescent (do not propagate). In contrast, AC driving leads to a non-zero transmission $\mathcal{T}_{RL}$ even for energies below $\Delta$, as evident from the green and blue curves in the figure. This is due to the fact that an incident electron from the left lead, possessing energy $E<\Delta$, can overcome the potential step and transfer to the right lead into a higher energy state ($E+n\hbar\omega>\Delta$) by absorbing $n$ energy quanta through AC driving. On the other hand, for $E<\Delta$, the red and magenta curves (right-to-left transmission, $\mathcal{T}_{LR}$) exhibit behavior analogous to the static case: they are zero because incident states from the right lead cannot propagate below the step. Figure 6(a) also shows that for energies above the step, all AC-driven curves lie slightly below their static counterparts (dashed lines).
@@
Similar to the findings in Section IV.1, albeit less pronounced, additional kink structures appear at $E=\Delta+\hbar\omega$ (marked by the vertical dotted lines). To enhance visibility of these features, in Fig. 6(a), we include an inset showing a zoom-in of the red and blue curves ($\Delta=0.75\mu$). Note that the kinks for the two transmissions in the blue and red curves go in opposite directions (one goes down and the other goes up, respectively), so that when summed together, they may even compensate. This is shown in Figure 6(b) where we plot the averaged transmissions $\bar{\mathcal{T}}=(\mathcal{T}_{RL}+\mathcal{T}_{LR})/2$. Note that here the range of values of transmissions is different from the one used in panel (a) and ranges between 0 and 0.9. Finally, both Figs. 6(a) and (b) show that there is no qualitative difference between the group of curves for $\Delta=1.1\mu$ and $\Delta=0.75\mu$, apart from the obvious shift in energy.
@@
Before proceeding, we note that because of the difference between $\mathcal{T}_{RL}(\epsilon)$ and $\mathcal{T}_{LR}(\epsilon)$, a non-zero pumping current can flow through the system even in the absence of voltage or thermal biases (i.e. $\Delta V=0$ and $\Delta T=0$). Indeed, according to Eqs. (3) and (5), in this case, the charge current is given by
$J_{e}=J_{e}^{P}=\frac{e}{h}\int_{-\infty}^{\infty}dE\left[\mathcal{T}_{LR}(E)-\mathcal{T}_{RL}(E)\right]f(E).$ (16)
@@
More details are reported in Appendix B. We can now study the thermoelectric properties of the system and calculate the coefficient $L_{\rm eh}$ defined in Eq. (8). Here we do not consider the Seebeck coefficient $S$ because the charge current, and hence the coefficient $L_{\rm ee}$, is strongly suppressed when the chemical potential $\mu$ is lower than the height of the step $\Delta$. Indeed, in this situation, the behavior of $S$ is totally dominated by $L_{\rm ee}$, see Eq. (6).
@@
Since in the static situation $L_{\rm eh}$ scales as $T^{3}$ in the low-temperature regime, see Sec. II, in Fig. 7 we plot the ratio $L_{\rm eh}/T^{3}$ as a function of the rescaled temperature $k_{B}T/\mu$ for the two values of potential step height: blue curve for $\Delta=0.75\mu$ and green curve for $\Delta=1.1\mu$. The dashed black curves refer for the static case. The main difference between the two is that, at $T=0$, one takes a finite value ($\Delta=0.75\mu$), while the other becomes zero ($\Delta=1.1\mu$). Both curves exhibit a broad peak at intermediate values of temperatures (between 0.05 and 0.1 $k_{B}T/\mu$). Figure 7(a) shows that the AC driving, in both cases, causes $L_{\rm eh}$ to slightly increase at small $T$ and slightly decrease at higher $T$. In both cases, however, the two curves do not deviate significantly from the static case. This behavior is discussed in detail in App. C.
@@
where $L_{\rm eh}^{(0)}$ denotes the static thermoelectric coefficient, i.e., the value of $L_{\rm eh}$ when $v=0$. The quantity $R_{L_{\rm eh}}$ is plotted in Fig. 7(b) as a function of the normalized temperature for the two different values of $\Delta$, for fixed values of $\omega$, $v$, and $\eta$. The blue and green curves represent the cases with the step value $\Delta=0.75\mu$ and $\Delta=1.1\mu$, respectively. Both curves show an enhancement of $L_{\rm eh}$ in the low-temperature regime. The blue curve has a maximum of about 5% increase, while the green curve shows a much larger increase, since it diverges for $T$ going to zero. Such a divergence is entirely due to the vanishing of the static coefficient $L_{\rm eh}^{(0)}$. Indeed, the photon-assisted effect of the driving enhances the thermoelectric coefficient $L_{\rm eh}$ in a range of temperatures where transport is otherwise suppressed by the presence of the potential step. The inset shows a logarithmic scale zoom-in in a range of small temperatures.
We stress that, from an experimental point of view, the changes due to the driving in $L_{\rm eh}$ are observable. Indeed, in a realistic nanowire setup, see for example Ref. [69], a chemical potential $\mu=1-10$ meV and a gap $\Delta=1.1\mu$ are achievable with gate-tunable devices. In this case, a change of $L_{\rm eh}$ on the order of $0.1(e/h)k_{B}^{2}/\mu$ [read out from the green curve in Fig. 7(a) at a small temperature] corresponds to detecting a current, see Eq. (5), $J_{\rm e}~\sim~2-20$ pA at $T=1$ K and $\Delta T=0.1$ K, which is within experimental reach.